# Carrier Aggregation Enabled MIMO-OFDM Integrated Sensing and Communication

### v1

@@ II-C2 Received sensing signal
At the receiver (Rx) of BS, the mixed echo signals on high and low-frequency bands are separated through matched filtering. Therefore, the received echo signal on the $n$-th subcarrier during the $m$-th OFDM symbol time is denoted by
$\mathbf{y}_{m,n}^{S}$ $=\mathbf{y}_{m,n}^{S,1}+\mathbf{y}_{m,n}^{S,2}$ (12)
$=\sum_{b=1}^{\mathcal{B}}\mathbf{H}_{m,n}^{b}\mathbf{x}_{m,n}^{b}+\mathbf{z}_{m,n}^{S},$
where $\mathbf{y}_{m,n}^{S}\in\mathbb{C}^{N_{\text{R}}\times 1}$ and $\mathbf{y}_{m,n}^{S,b}$ denotes the echo signal in $b$-th CC; $\mathbf{x}_{m,n}^{b}\in\mathbb{C}^{N_{\text{T}}\times 1}=\mathbf{X}_{n}^{b}\Upsilon_{m}$ is the transmit data with $\Upsilon_{m}=[0_{0},0_{1},\cdots,1_{m},\cdots,0_{M_{b}-1}]^{\text{T}}\in\mathbb{C}^{M_{b}\times 1}$; $\mathbf{z}_{m,n}^{S}\sim\mathcal{CN}(0,\sigma_{S}^{2})\in\mathbb{C}^{N_{\text{R}}\times 1}$ is an AWGN vector.
@@ III A Novel CA-enabled MIMO-OFDM ISAC Signal Processing Method
In this section, a novel CA-enabled MIMO-OFDM ISAC signal processing method is proposed, as shown in Fig. 2. Without loss of generality, we assume that the $1$-st CC is on the low-frequency band and the $2$-nd CC is on the high-frequency band. Through matched filtering at the Rx, the mixed echo signals on the two CCs are separated.
There are two challenges in sensing signal processing in the proposed ISAC system: 1) The initial phase misalignment of the echo signals on high and low-frequency bands due to attenuation and RCS; 2) The fusion of the sensing data on high and low-frequency bands with different physical-layer parameters, such as subcarrier spacing and symbol duration. Therefore, we consider both data-level and symbol-level fusion methods of the sensing data on high and low-frequency bands.
The data-level fusion method is as follows: After the separation of mixed echo signals on high and low-frequency bands, the joint angle-range-velocity estimation method proposed in [21] is applied to the echo signal on high-frequency band to yield the estimated range $r_{\text{data}}^{1}$ and velocity $v_{\text{data}}^{1}$ of target. Similarly, the estimated range $r_{\text{data}}^{2}$ and velocity $v_{\text{data}}^{2}$ of target are obtained by processing the echo signal on low-frequency band. Finally, the final estimation of range is $\hat{r}_{\text{data}}=(r_{\text{data}}^{1}+r_{\text{data}}^{2})/2$, while the final estimation of velocity is $\hat{v}_{\text{data}}=(v_{\text{data}}^{1}+v_{\text{data}}^{2})/2$. However, the sensing accuracy of data-level fusion method is low. Therefore, this paper proposes a two-stage symbol-level fusion method, including signal preprocessing stage and sensing information fusion stage, detailed in Section III-A.
@@ III-A Signal Preprocessing Stage
Facing the above challenges, one of the purposes of sensing preprocessing is to align the initial phases of the echo signals on high and low-frequency bands.
@@ III-A1 Angle estimation
With identical AoA and AoD for two CCs and high signal-to-noise ratio (SNR) of the echo signal on low-frequency band, we utilize the echo signal on low-frequency band to estimate AoA and AoD.
Firstly, we remove the transmit data $\mathbf{x}_{m,n}^{1}$ in $\mathbf{y}_{m,n}^{S,1}$ using known communication data, and the echo signal on low-frequency band by all antennas is
$\hat{\mathbf{H}}_{m,n}^{1}$ $=\mathbf{y}_{m,n}^{S,1}\left(\mathbf{x}_{m,n}^{1}\right)^{\text{H}}\left(\mathbf{x}_{m,n}^{1}\left(\mathbf{x}_{m,n}^{1}\right)^{\text{H}}\right)^{-1}$ (13)
$=\kappa_{S}^{b}e^{j2\pi\left(f_{\text{s}}^{b}mT^{b}-n\Delta f^{b}\tau_{0}\right)}\mathbf{a}_{\text{Rx}}(\theta_{\text{Rx}})\mathbf{a}_{\text{Tx}}^{\text{T}}(\theta_{\text{Tx}})\mathbf{I}+\mathbf{z}_{m,n}^{S,1},$
where $\hat{\mathbf{H}}_{m,n}^{1}\in\mathbb{C}^{N_{\text{R}}\times N_{\text{T}}}$ and $\mathbf{I}\in\mathbb{C}^{N_{\text{T}}\times N_{\text{T}}}$ is the identity matrix; $\mathbf{z}_{m,n}^{S,1}$ is an AWGN matrix.
Upon observing $\hat{\mathbf{H}}_{m,n}^{1}$, it is obvious that the AoA introduces a linear phase shift along the received antenna elements. The AoA estimation is obtained using the conventional multiple signal classification (MUSIC) method [22], with the following steps.
Step 1: Calculate the covariance matrix $\mathbf{R}_{\hat{\mathbf{H}}_{m,n}^{1}}$ of signal $\hat{\mathbf{H}}_{m,n}^{1}$.
$\mathbf{R}_{\hat{\mathbf{H}}_{m,n}^{1}}=\hat{\mathbf{H}}_{m,n}^{1}\left[\hat{\mathbf{H}}_{m,n}^{1}\right]^{\text{H}}.$ (14)
Step 2: $\mathbf{R}_{\hat{\mathbf{H}}_{m,n}^{1}}$ is performed an eigenvalue decomposition (EVD) to obtain
$\text{eig}\left(\mathbf{R}_{\hat{\mathbf{H}}_{m,n}^{1}}\right)=\mathbf{U}_{\text{s}}\mathbf{\Lambda}_{\text{s}}\mathbf{U}_{\text{s}}^{\text{H}}+\mathbf{U}_{\text{n}}\mathbf{\Lambda}_{\text{n}}\mathbf{U}_{\text{n}}^{\text{H}},$ (15)
where $\mathbf{\Lambda}_{\text{s}}$ and $\mathbf{\Lambda}_{\text{n}}$ denote the diagonal matrix of signal and noise, respectively; $\mathbf{U}_{\text{s}}$ and $\mathbf{U}_{\text{n}}$ represent the subspaces of signal and noise, respectively. The [22] proves that the noise subspace is orthogonal to the signal subspace. Therefore, we traverse the signal subspace to obtain AoA estimation.
@@
Step 4: Search the peak of $f_{\text{music}}$ to obtain the AoA estimation $\hat{\theta}_{\text{Rx}}$.
With the MUSIC method, the estimation result $\hat{\theta}_{\text{Rx}}$ of AoA is acquired for phase compensation.
III-A2 Phase compensation
@@
Phase compensation is designed to cancel known communication data, AoA, and AoD information contained in the echo signals on high and low-frequency bands. The echo signal by all antennas on the $n$-th subcarrier during the $m$-th OFDM symbol time in the $b$-th CC is expressed as
@@
$\mathbf{y}_{m,n}^{S,b}=\kappa_{S}^{b}e^{j2\pi\left(f_{\text{s}}^{b}mT^{b}-n\Delta f^{b}\tau_{0}\right)}\mathbf{a}_{\text{Rx}}(\theta_{\text{Rx}})\mathcal{X}_{m,n}^{b}+\mathbf{z}_{m,n}^{S,b},$ (18)
where $\mathbf{y}_{m,n}^{S,b}\in\mathbb{C}^{N_{\text{R}}\times 1}$ and $\mathcal{X}_{m,n}^{b}=\mathbf{a}_{\text{Tx}}^{\text{T}}(\theta_{\text{Tx}})\mathbf{x}_{m,n}^{b}\in\mathbb{C}$ is a transmit factor.
With the known transmit data and AoD estimation, we can compensate for $\mathcal{X}_{m,n}^{b}$ by multiplying the compensation factor $\widetilde{\mathcal{X}}_{m,n}^{b}=\left(\mathbf{a}_{\text{Tx}}^{\text{T}}(\hat{\theta}_{\text{Tx}})\mathbf{x}_{m,n}^{b}\right)^{*}$ on both sides of (18), denoted by
$\mathbf{y}_{m,n}^{S,b}\widetilde{\mathcal{X}}_{m,n}^{b}=$ (19)
@@
$\hat{\mathcal{X}}_{m,n}^{b}\kappa_{S}^{b}e^{j2\pi\left(f_{\text{s}}^{b}mT^{b}-n\Delta f^{b}\tau_{0}\right)}\mathbf{a}_{\text{Rx}}(\theta_{\text{Rx}})+\mathbf{z}_{m,n}^{S,b}\widetilde{\mathcal{X}}_{m,n}^{b},$
where $\hat{\mathcal{X}}_{m,n}^{b}=\mathcal{X}_{m,n}^{b}\widetilde{\mathcal{X}}_{m,n}^{b}\in\mathbb{R}$. Then, with the AoA estimation, $\mathbf{a}_{\text{Rx}}(\theta_{\text{Rx}})$ in (19) is compensated, and $\mathbf{r}_{m,n}^{S,b}$ is expressed as
$\mathbf{r}_{m,n}^{S,b}=\mathbf{h}_{m,n}^{S,b}+\mathbf{n}_{m,n}^{S,b},$ (20)
where
$\mathbf{r}_{m,n}^{S,b}\in\mathbb{C}^{N_{\text{R}}\times 1}=\mathbf{y}_{m,n}^{S,b}\widetilde{\mathcal{X}}_{m,n}^{b}\circ\mathbf{a}_{\text{Rx}}^{*}(\theta_{\text{Rx}}),$ (21)
$\mathbf{n}_{m,n}^{S,b}=\mathbf{z}_{m,n}^{S,b}\widetilde{\mathcal{X}}_{m,n}^{b}\circ\mathbf{a}_{\text{Rx}}^{*}(\theta_{\text{Rx}}),$ (22)
$\mathbf{h}_{m,n}^{S,b}=\hat{\mathcal{X}}_{m,n}^{b}\kappa_{S}^{b}e^{j2\pi\left(f_{\text{s}}^{b}mT^{b}-n\Delta f^{b}\tau_{0}\right)}\mathbf{a}_{\text{Rx}}(\theta_{\text{Rx}})\circ\mathbf{a}_{\text{Rx}}^{*}(\theta_{\text{Rx}}).$ (23)
III-A3 Space-domain diversity
Upon observing Eq. (23), the echo signals on each antenna are identical and can be coherently accumulated to improve SNR. The echo signal after coherent accumulation on the $n$-th subcarrier during the $m$-th OFDM symbol time in the $b$-th CC is expressed as
$\mathbf{d}_{m,n}^{S,b}=\frac{1}{N_{\text{R}}}\sum_{p=0}^{N_{\text{R}}-1}\mathbf{r}_{m,n}^{S,b}(p),$ (24)
@@ III-A3 Space-domain diversity
where $\mathbf{r}_{m,n}^{S,b}(p)$ is the $p$-th element of $\mathbf{r}_{m,n}^{S,b}$. According to (24), the echo signal on $N_{b}$ subcarriers during $M_{b}$ OFDM symbol times in the $b$-th CC is written in matrix form as
$\mathbf{D}^{S,b}=\mathbf{S}^{S,b}+\mathbf{Z}^{S,b},$ (25)
@@
where $\mathbf{Z}^{S,b}$ is an AWGN matrix and $\mathbf{S}^{S,b}\in\mathbb{C}^{N_{b}\times M_{b}}$ denotes a delay-Doppler information matrix, expressed in (26).
@@
$\centering\mathbf{S}^{S,b}=\hat{\mathcal{X}}_{m,n}^{b}\kappa_{S}^{b}\left[\begin{array}[]{cccc}1&e^{j2\pi f_{\text{s}}^{b}T^{b}}&\cdots&e^{j2\pi(M_{b}-1)f_{\text{s}}^{b}T^{b}}\\ e^{-j2\pi\Delta f^{b}\tau_{0}}&e^{-j2\pi\Delta f^{b}\tau_{0}}e^{j2\pi f_{\text{s}}^{b}T^{b}}&\cdots&e^{-j2\pi\Delta f^{b}\tau_{0}}e^{j2\pi(M_{b}-1)f_{\text{s}}^{b}T^{b}}\\ \vdots&\vdots&\ddots&\vdots\\ e^{-j2\pi(N_{b}-1)\Delta f^{b}\tau_{0}}&e^{-j2\pi(N_{b}-1)\Delta f^{b}\tau_{0}}e^{j2\pi f_{\text{s}}^{b}T^{b}}&\cdots&e^{-j2\pi(N_{b}-1)\Delta f^{b}\tau_{0}}e^{j2\pi(M_{b}-1)f_{\text{s}}^{b}T^{b}}\\ \end{array}\right].\@add@centering$ (26)
III-A4 CCC method

### v2

@@ II-C2 Received sensing signal
[33] verifies that the high and low-frequency band signals have similar behavior, that is, they reach the receiver (Rx) of BS through the same line of sight path and target reflection. At the Rx of BS, the mixed sensing echo signals on high and low-frequency bands are separated through matched filtering. Therefore, the received echo signal on the $n$-th subcarrier during the $m$-th OFDM symbol time is denoted by
$\mathbf{y}_{m,n}^{S}=\sum_{b=1}^{\mathcal{B}}\mathbf{y}_{m,n}^{S,b}=\sum_{b=1}^{\mathcal{B}}\mathbf{H}_{m,n}^{b}\mathbf{W}_{\text{Tx}}^{b}\mathbf{d}_{m,n}^{b}+\mathbf{z}_{m,n}^{S,b},$ (12)
where $\mathbf{y}_{m,n}^{S}\in\mathbb{C}^{N_{\text{R}}\times 1}$ and $\mathbf{y}_{m,n}^{S,b}$ denotes the echo signal in $b$-th CC; $\mathbf{W}_{\text{Tx}}^{b}\in\mathbb{C}^{N_{\text{T}}\times N_{\text{T}}}$ represents the transmit beamforming matrix used to point in the direction of interest, and $\mathbf{d}_{m,n}^{b}\in\mathbb{C}^{N_{\text{T}}\times 1}$ is the dedicated sensing data with constant mode 1; $\mathbf{z}_{m,n}^{S,b}\sim\mathcal{CN}(0,\sigma_{S,b}^{2})\in\mathbb{C}^{N_{\text{R}}\times 1}$ is an AWGN vector, which characterizes the different noise variances of high and low-frequency bands.
@@ III A Novel CA-enabled MIMO-OFDM ISAC Signal Processing Method
In this section, a novel CA-enabled MIMO-OFDM ISAC signal processing method is proposed, as illustrated in Fig. 2. Without loss of generality, we assume that the $1$-st CC is on the low-frequency band and the $2$-nd CC is on the high-frequency band. Through matched filtering at the Rx, the mixed sensing echo signals on the two CCs are separated.
There are two challenges in sensing signal processing within the proposed ISAC system: 1) The initial phase misalignment of the echo signals on high and low-frequency bands due to attenuation and RCS; 2) The fusion of the sensing data on high and low-frequency bands with different physical-layer parameters, such as subcarrier spacing and symbol duration. Therefore, we consider both data-level and symbol-level fusion methods of the sensing data on high and low-frequency bands.
The data-level fusion method is as follows: After the separation of mixed echo signals on high and low-frequency bands, the joint angle-range-velocity estimation method proposed in [34] is applied to the echo signal on high-frequency band to yield the estimated range set $\{r_{i,\text{data}}^{1}\}_{i=1,2,\cdots,I}$ and velocity set $\{v_{i,\text{data}}^{1}\}_{i=1,2,\cdots,I}$ of $I$ targets, which have a one-to-one relationship. Similarly, the estimated range set $\{r_{i,\text{data}}^{2}\}_{i=1,2,\cdots,I}$ and velocity set $\{v_{i,\text{data}}^{2}\}_{i=1,2,\cdots,I}$ of $I$ targets are obtained by processing the echo signal on low-frequency band. We combine the difference of accuracy and SNR to give the data-level fusion weights, and the variances of received signal on high and low-frequency bands can be obtained by some methods, such as maximum likelihood estimation (MLE) [35]. Therefore, the final estimation for range of the $i$-th target is $\hat{r}_{i,\text{data}}=r_{i,\text{data}}^{1}+\frac{\sigma_{S,1}^{2}\frac{\Delta f^{2}M_{b}}{\Delta f^{1}}}{\sigma_{S,1}^{2}\frac{\Delta f^{2}M_{b}}{\Delta f^{1}}+\sigma_{S,2}^{2}}(r_{i,\text{data}}^{2}-r_{i,\text{data}}^{1})$, while the final estimation for velocity of the $i$-th target is $\hat{v}_{i,\text{data}}=v_{i,\text{data}}^{1}+\frac{\sigma_{S,1}^{2}\frac{\Delta f^{2}N_{b}}{\Delta f^{1}}}{\sigma_{S,1}^{2}\frac{\Delta f^{2}N_{b}}{\Delta f^{1}}+\sigma_{S,2}^{2}}(v_{i,\text{data}}^{2}-v_{i,\text{data}}^{1})$ [28]. However, the sensing accuracy of data-level fusion method is low. Therefore, this paper proposes a two-stage symbol-level fusion method, including signal preprocessing stage and sensing information fusion stage, detailed in Section III-A.
@@ III-A Signal Preprocessing Stage
One primary purposes of sensing preprocessing is to align the initial phases of the echo signals on high and low-frequency bands.
@@ III-A1 Angle estimation
With identical AoA and AoD for two CCs and high signal-to-noise ratio (SNR) of the echo signal on low-frequency band, we utilize the echo signal on low-frequency band to estimate AoA.
Firstly, we remove the known transmit data $\mathbf{d}_{m,n}^{1}$ in $\mathbf{y}_{m,n}^{S,1}$ by right-multiplying the generalized inverse matrix $\mathbf{X}_{m,n}^{\dagger}=\left(\mathbf{d}_{m,n}^{1}\right)^{\text{H}}\left(\mathbf{d}_{m,n}^{1}\left(\mathbf{d}_{m,n}^{1}\right)^{\text{H}}+\varrho\mathbf{I}\right)^{-1}$, and the echo signal on low-frequency band by all antennas is
$\hat{\mathbf{H}}_{m,n}^{1}$ $=\mathbf{y}_{m,n}^{S,1}\mathbf{X}_{m,n}^{\dagger}$ (13)
$\approx\sum_{i=1}^{I}\left[\begin{array}[]{l}\kappa_{i,S}^{1}e^{j2\pi\left(f_{i,\text{s}}^{1}mT^{1}-n\Delta f^{1}\tau_{i,0}\right)}\\ \mathbf{a}_{\text{Rx}}(\theta_{i,\text{Rx}})\mathbf{a}_{\text{Tx}}^{\text{T}}(\theta_{i,\text{Tx}})\mathbf{I}\end{array}\right]\mathbf{W}_{\text{Tx}}^{b}+\mathbf{z}_{m,n}^{S,1}\mathbf{X}_{m,n}^{\dagger},$
where $\hat{\mathbf{H}}_{m,n}^{1}\in\mathbb{C}^{N_{\text{R}}\times N_{\text{T}}}$ and $\mathbf{I}\in\mathbb{C}^{N_{\text{T}}\times N_{\text{T}}}$ is the identity matrix. $\varrho$ denotes a regularization parameter introduced to improve the accuracy of the numerical recovery [19].
Upon observing $\hat{\mathbf{H}}_{m,n}^{1}$, it is obvious that the AoA introduces a linear phase shift along the receive antenna elements. The AoA estimations can be obtained by multiple signal classification (MUSIC) method, estimating signal parameter via rotational invariance techniques method, and the advanced methods based on complex neural network [3, 36]. Due to its ability to achieve parameter estimation in non-contiguous signals in the space-time-frequency-code domains with high precision, we take the MUSIC method as an example [37, 28], with the following steps.
Step 1: Calculate the covariance matrix $\mathbf{R}_{\hat{\mathbf{H}}_{m,n}^{1}}$ of $\hat{\mathbf{H}}_{m,n}^{1}$.
$\mathbf{R}_{\hat{\mathbf{H}}_{m,n}^{1}}=\frac{\hat{\mathbf{H}}_{m,n}^{1}\left[\hat{\mathbf{H}}_{m,n}^{1}\right]^{\text{H}}}{N_{1}M_{1}}\in\mathbb{C}^{N_{\text{R}}\times N_{\text{R}}}.$ (14)
Step 2: $\mathbf{R}_{\hat{\mathbf{H}}_{m,n}^{1}}$ is performed an eigenvalue decomposition $\text{eig}(\cdot)$ to obtain
$\text{eig}\left(\mathbf{R}_{\hat{\mathbf{H}}_{m,n}^{1}}\right)=\left[\mathbf{U}_{\text{s}},\mathbf{\Lambda}_{\text{s}}\right],$ (15)
where $\mathbf{\Lambda}_{\text{s}}$ denotes a diagonal matrix with descending order of eigenvalues. $\mathbf{U}_{\text{s}}$ is the orthogonal eigenmatrix. The number of incident signals can be accurately ascertained by estimating the rank of the signal space from the differential vector derived from the eigenvalues of the signal matrix [38], which are assumed to be the same as the number of targets. Then, the noise subspace $\mathbf{U}_{\text{n}}=\left[\mathbf{U}_{\text{s}}\right]_{:,I+1:N_{\text{R}}}$ is obtained, which is used to calculate the spatial spectral function.
@@
Step 4: Search the $I$ peaks of $f_{\text{music}}$ to obtain the AoA estimations set $\{\hat{\theta}_{i^{\prime},\text{Rx}}\}_{i^{\prime}=1,2\cdots,I}$.
III-A2 Spatial filtering
To mitigate target ghosts introduced by potential nonlinear processing, we employ spatial filtering operation to separate the echo signals that are blended from multiple targets.
@@ III-A2 Spatial filtering
Similar to (13), after the known transmit data $\mathbf{d}_{m,n}^{b}$ is removed, the echo signal by all antennas on the $n$-th subcarrier during the $m$-th OFDM symbol time in the $b$-th CC is expressed as
@@
$\hat{\mathbf{H}}_{m,n}^{b}\approx\sum_{i=1}^{I}\left[\begin{array}[]{l}\kappa_{i,S}^{b}e^{j2\pi\left(f_{i,\text{s}}^{b}mT^{b}-n\Delta f^{b}\tau_{i,0}\right)}\\ \mathbf{a}_{\text{Rx}}(\theta_{i,\text{Rx}})\mathbf{a}_{\text{Tx}}^{\text{T}}(\theta_{i,\text{Tx}})\mathbf{I}\end{array}\right]\mathbf{W}_{\text{Tx}}^{b}+\mathbf{Z}_{m,n}^{S,b},$ (18)
where $\mathbf{Z}_{m,n}^{S,b}$ is a noise matrix.
Similar to the principle of OFDM signal demodulation, we can separate the individual echo signal of each target from $\hat{\mathbf{H}}_{m,n}^{b}$ with the estimated AoAs $\{\hat{\theta}_{i^{\prime},\text{Rx}}\}_{i^{\prime}=1,2\cdots,I}$. The echo signal of the $i^{\prime}$-th target by all antennas on the $n$-th subcarrier during the $m$-th OFDM symbol time in the $b$-th CC is expressed in (19). When the target angle difference is greater than the angular resolution of the antenna, and the SNR is not poor, the magnitude of the Term 1 is much greater than that of the Term 2. Fortunately, the spatial filtering operation can confer substantial SNR gains in space-domain.
$\hat{\mathbf{H}}_{m,n}^{b,i^{\prime}}$ $=\mathbf{a}_{\text{Rx}}^{\text{H}}(\hat{\theta}_{i^{\prime},\text{Rx}})\hat{\mathbf{H}}_{m,n}^{b}(\mathbf{W}_{\text{Tx}}^{b})^{\text{H}}\mathbf{a}_{\text{Tx}}^{\text{*}}(\hat{\theta}_{i^{\prime},\text{Tx}})/(N_{\text{R}}N_{\text{T}})$ (19)
$\approx\underbrace{\kappa_{i^{\prime},S}^{b}e^{j2\pi\left(f_{i^{\prime},\text{s}}^{b}mT^{b}-n\Delta f^{b}\tau_{i^{\prime},0}\right)}}_{\text{The}\ i^{\prime}\text{-th target}\ (\textbf{Term 1})}$
@@
$\quad+\underbrace{\mathbf{a}_{\text{Rx}}^{\text{H}}(\hat{\theta}_{i^{\prime},\text{Rx}})\left[\sum_{i=1,i\neq i^{\prime}}^{I}\left[\begin{array}[]{l}\kappa_{i,S}^{b}e^{j2\pi\left(f_{i,\text{s}}^{b}mT^{b}-n\Delta f^{b}\tau_{i,0}\right)}\\ \mathbf{a}_{\text{Rx}}(\theta_{i,\text{Rx}})\mathbf{a}_{\text{Tx}}^{\text{T}}(\theta_{i,\text{Tx}})\mathbf{I}\end{array}\right]\mathbf{W}_{\text{Tx}}^{b}+\mathbf{Z}_{m,n}^{S,b}\right](\mathbf{W}_{\text{Tx}}^{b})^{\text{H}}\mathbf{a}_{\text{Tx}}^{\text{*}}(\hat{\theta}_{i^{\prime},\text{Tx}})/(N_{\text{R}}N_{\text{T}}).}_{\text{Interference}\ (\textbf{Term 2})}$
@@
According to (19), the echo signal of the $i^{\prime}$-th target on $N_{b}$ subcarriers during $M_{b}$ OFDM symbol times in the $b$-th CC is rewritten in matrix form as
$\mathbf{D}_{i^{\prime}}^{S,b}=\mathbf{S}_{i^{\prime}}^{S,b}+\mathbf{Z}_{i^{\prime}}^{S,b},$ (20)
@@
where $\mathbf{Z}_{i^{\prime}}^{S,b}$ is an interference matrix including the interference from other target and noise, and $\mathbf{S}_{i^{\prime}}^{S,b}\in\mathbb{C}^{N_{b}\times M_{b}}$ denotes a delay-Doppler information matrix, expressed in (21).
@@
$\centering\mathbf{S}_{i^{\prime}}^{S,b}=\kappa_{i^{\prime},S}^{b}\left[\begin{array}[]{cccc}1&e^{j2\pi f_{i^{\prime},\text{s}}^{b}T^{b}}&\cdots&e^{j2\pi(M_{b}-1)f_{i^{\prime},\text{s}}^{b}T^{b}}\\ e^{-j2\pi\Delta f^{b}\tau_{i^{\prime},0}}&e^{-j2\pi\Delta f^{b}\tau_{i^{\prime},0}}e^{j2\pi f_{i^{\prime},\text{s}}^{b}T^{b}}&\cdots&e^{-j2\pi\Delta f^{b}\tau_{i^{\prime},0}}e^{j2\pi(M_{b}-1)f_{i^{\prime},\text{s}}^{b}T^{b}}\\ \vdots&\vdots&\ddots&\vdots\\ e^{-j2\pi(N_{b}-1)\Delta f^{b}\tau_{i^{\prime},0}}&e^{-j2\pi(N_{b}-1)\Delta f^{b}\tau_{i^{\prime},0}}e^{j2\pi f_{i^{\prime},\text{s}}^{b}T^{b}}&\cdots&e^{-j2\pi(N_{b}-1)\Delta f^{b}\tau_{i^{\prime},0}}e^{j2\pi(M_{b}-1)f_{i^{\prime},\text{s}}^{b}T^{b}}\\ \end{array}\right].\@add@centering$ (21)
III-A3 CCC method